# Motivic interpretations for iterated integrals on some specific algebraic curves

### v2

@@ 1. Introduction
This conjecture has not yet been proved as of March 2024, but Goncharov ([7]), Terasoma ([15]), Deligne–Goncharov ([4]) showed
@@ 1.1. Main Results
The main result of this paper is to give a motivic interpretation for these values by using motivic iterated integrals. To explain the claim, we need more notation. Fix an algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$. For a number field $F(\subseteq\overline{\mathbb{Q}})$ and a $\mathbb{Q}$-vector subspace $R\subseteq F^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$, let $\mathcal{A}_{F,R}$ be the fundamental Hopf algebra of $R$ and $\mathcal{H}_{F,R}:=\mathcal{A}_{F,R}\otimes_{\mathbb{Q}}\mathbb{Q}[\tau]$. Let $\mathrm{per}:\mathcal{H}_{F,R}\rightarrow\mathbb{C}$ be the period map. Then the Galois group $\mathrm{Gal}(F/\mathbb{Q})$ acts on $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ (see Section 4 for details).
@@
where $\xi_{N}:=\exp\left(\frac{2\pi\sqrt{-1}}{N}\right)$. We introduce a $\mathbb{Q}$-vector subspace $R_{f}\subseteq F_{f}^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ which is defined by
$R_{f}:=\begin{cases}\langle 2\rangle_{\mathbb{Q}},&f=g,\\ \langle 2,3\rangle_{\mathbb{Q}},&f=h.\end{cases}$
Also, we put $G_{f}:=\mathrm{Gal}(F_{f}/\mathbb{Q})$. Then, we have the following theorem.
@@
We will give the value of $D_{f}^{(k)}$ and conjectural value of $\dim_{\mathbb{Q}}$ after Corollary 5.10 obtained by numerical experiments.
@@
We organize this paper as follows. In Section 2, we set up some notations and definitions. In Section 3, we explain some basic properties of our iterated integrals and give some explicit formulas. These formulas shows our periods can be written as linear combinations of well known periods and it reads us to seek their motivic interpretations, which are the contents in Sections 4 and 5. In Section 4, we summarize the basic properties of the motivic iterated integrals given by Deligne, Glanois, and Goncharov to give motivic interpretations of the iterated integrals on our algebraic curves. In Section 5, we give motivic interpretations and investigate their algebraic structures. This interpretation explains the appearance of $\pi$, $\log 2$, $\log 3$, $L(k,\chi_{-3})$, and Riemann zeta values in the explicit formulas in Example 3.1-(3), Propositions 3.5 and 3.7. We also notice about the direct sum decomposition of the $\mathbb{Q}$-linear space $Z_{f}^{(k)}$. In Section 5.4, we describe the relationship between our periods and special values called alternating multiple mixed values, defined by Xu–Yan–Zhao [16] and further studied by Charlton [1].
We remark that we considered in [14] only the values given by iterated integrals with 33 33 elements of $B_{g}$ and we consider a more wider class in this paper.
@@ Acknowledgement
The author would like to express his deepest gratitude to his supervisor, Professor T. Yamauchi. In particular, we follow his idea in a part of the computation in Section 3. The author would like to thank Professor H. Furusho of Nagoya University for inviting me to his monthly seminar and giving me various advice of my research. The author would also like to thank Professor M. Kaneko of Kyushu University for inviting me to the multiple zeta seminar and discussing multiple $\widetilde{T}$-values. The author is grateful to S. Charlton of the Max Planck Institute for Mathematics for discussing alternating multiple mixed values. The author would like to thank the WISE Program for AI Electrics of Tohoku University for their financial support to carry out our study.
@@ 2. Preparations
$\displaystyle L\left(\begin{matrix}k_{1},\dots,k_{d}\\ \alpha_{1},\dots,\alpha_{d}\end{matrix}\right):=(-1)^{d}\mathrm{I}_{\mathrm{dch}}(0;a_{1}^{-1},\{0\}^{k_{1}-1},\dots,\alpha_{d}^{-1},\{0\}^{k_{d}-1};1)$ (19)
@@
If $f(X,Y,Z)=X^{2}+Y^{2}-Z^{2}$, we can take a $\mathbb{Q}$-basis as $B_{f}=\left\{\frac{\leavevmode\nobreak\ dx}{x},\frac{\leavevmode\nobreak\ dx}{1-x},\frac{\leavevmode\nobreak\ dx}{y},\frac{\leavevmode\nobreak\ dx}{xy}\right\}$. Then, if we restrict the 11 11-forms to the subset $\left\{\frac{\leavevmode\nobreak\ dx}{x},\frac{\leavevmode\nobreak\ dx}{1-x},\frac{\leavevmode\nobreak\ dx}{y}\right\}\subseteq B_{f}$, the iterated integrals equation 9 give $F_{2}$-MZVs considered in [14].
@@
    $\displaystyle\phi_{g}:\mathbb{P}^{1}(\mathbb{C})\overset{\sim}{\longrightarrow}X_{g},\phi_{g}(\lambda):=(x,y)=\left(\frac{2\lambda}{1+\lambda^{2}},\frac{1-\lambda^{2}}{1+\lambda^{2}}\right).$ (20)
    The points on $X_{g}$ with $x=0,1,\infty$ correspond to the following 55 55 points: $(x,y)=\left(0,\pm 1\right),(1,0)$ and $[X:Y:Z]=\left[1:\pm\sqrt{-1}:0\right]$. The dimension of $H^{1}_{\mathrm{dR}}(Y_{g}/\mathbb{Q})$ is 44 44 and we can take $B_{g}$ as
@@
    $\displaystyle\phi_{h}:\mathbb{P}^{1}(\mathbb{C})\overset{\sim}{\longrightarrow}X_{h},\phi_{h}(\lambda):=(x,y)=\left(\frac{\lambda(\lambda+2)}{1+\lambda+\lambda^{2}},\frac{1-\lambda^{2}}{1+\lambda+\lambda^{2}}\right).$ (21)
    The points on $X_{h}$ with $x=0,1,\infty$ correspond to the following 66 66 points: $(x,y)=(0,\pm 1),(1,0),(1,-1)$ and $[X:Y:Z]=[1:\xi_{3}^{\pm 1}:0]$. The dimension of $H^{1}_{\mathrm{dR}}(Y_{h}/\mathbb{Q})$ is 55 55 and we can take $B_{h}$ as
@@ 3.2. Explicit formula
where $\varphi_{j}=\sum_{n=0}^{\infty}c(n,\varphi_{j})x^{n-1}\leavevmode\nobreak\ dx$ (see [14, p.8, Proposition 3.1.3]). It follows equation 33 since $\omega_{s}=\sum_{n=0}^{\infty}c(n,\omega_{s})x^{n-1}\leavevmode\nobreak\ dx$ for each $s=0,1,\dots,6$. ∎
@@
To prove this explicit formula, we read the following lemma.
@@
If $z_{0}\in\mathbb{C}$ satisfies $|z_{0}|=1$ and $-\frac{\pi}{2}\leq\arg(z_{0})\leq\frac{\pi}{2}$, they hold
@@
$\displaystyle\mathrm{Li}_{k}(z_{0}):=-\mathrm{I}_{\mathrm{dch}_{0,z}}(0;1,\{0\}^{k-1};z_{0})=\sum_{n=1}^{\infty}\frac{z_{0}^{n}}{n^{k}}.$ (47)
@@
where $\gamma_{1}=\mathrm{dch}_{0,1-\varepsilon}$, $\gamma_{2}=C_{1}(1-\varepsilon,1+\varepsilon\sqrt{-1})$, and $\gamma_{3}=C_{0}(1+\varepsilon\sqrt{-1},z_{2})$ for sufficiently small $\varepsilon>0$ (see FIGURE 1).
@@
Рис. 1. integral paths
By $\varepsilon\searrow 0$ and calculating the regularization of the iterated integral (Remark 2.2), they hold
$\displaystyle\int_{\gamma_{1}}\frac{\leavevmode\nobreak\ dz_{1}}{1-z_{1}}=0$ (50)
@@
where $\gamma_{3}=\mathrm{dch}$ and $\gamma_{4}=C_{0}(1,z_{2})$. Here,
$\displaystyle\int_{\gamma}\frac{\leavevmode\nobreak\ dz_{1}}{1+z_{1}}=\log 2.$ (55)
@@
$\displaystyle=-\int_{\mathrm{dch}_{-1,-z_{0}}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}$ (57)
$\displaystyle=-\int_{\gamma_{5}\gamma_{6}\gamma_{7}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z},$ (58)
where $\gamma_{5}=\mathrm{dch}_{-1,-\varepsilon}$, $\gamma_{6}=C_{0}(-\varepsilon,\varepsilon\arg(-z_{0}))$, and $\gamma_{7}=\mathrm{dch}_{\varepsilon\arg(-z_{0}),-z_{0}}$ for sufficiently small $\varepsilon>0$ (see FIGURE 2).
@@
$\displaystyle\int_{\gamma_{7}}\frac{\leavevmode\nobreak\ dz}{z}=\int_{\gamma_{6}}\frac{\leavevmode\nobreak\ dz}{1-z}=0,$ (59)
@@
$\displaystyle\int_{\gamma_{5}\gamma_{6}\gamma_{7}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}=\int_{\gamma_{5}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}+\int_{\gamma_{5}}\frac{\leavevmode\nobreak\ dz}{1-z}\int_{\gamma_{6}}\frac{\leavevmode\nobreak\ dz}{z}+\int_{\gamma_{7}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}.$ (60)
Here, each terms can be calculated as
$\displaystyle\int_{\gamma_{5}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}$ $\displaystyle=\int_{\gamma_{5}^{-1}}\frac{\leavevmode\nobreak\ dz}{z}\frac{\leavevmode\nobreak\ dz}{1-z}$ (61)
@@
$\displaystyle\int_{\gamma_{5}}\frac{\leavevmode\nobreak\ dz}{1-z}\int_{\gamma_{6}}$ $\displaystyle\frac{\leavevmode\nobreak\ dz}{z}=\log 2\log z_{0},$ (64)
$\displaystyle\int_{\gamma_{7}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}$ $\displaystyle=\int_{\mathrm{dch}_{0,-z_{0}}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}=\mathrm{Li}_{2}(-z_{0}),$ (65)
@@
where $z=C_{0}(1,\sqrt{-1})$. Since $0\leq\arg(z)\leq\frac{\pi}{2}$ for any $z\in\gamma([0,1])$, we have
@@
$\displaystyle=\left[\log^{k-2}z(-\mathrm{Li}_{2}(z)+\frac{\pi\sqrt{-1}}{2}\log z-\mathrm{Li}_{2}(-z)+\log 2\log z+\frac{1}{2}\log^{2}z)\right]_{1}^{\sqrt{-1}}$ (74)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -(k-2)\int_{1}^{\sqrt{-1}}\log^{k-3}z(-\mathrm{Li}_{2}(z)+\frac{\pi\sqrt{-1}}{2}\log z-\mathrm{Li}_{2}(-z)+\log 2\log z+\frac{1}{2}\log^{2}z)\frac{\leavevmode\nobreak\ dz}{z}.$ (75)
@@
$\displaystyle=\sum_{j=2}^{k}(-1)^{j-1}\frac{(k-2)!}{(k-j)!}\left(\frac{\pi\sqrt{-1}}{2}\right)^{k-j}\left(\mathrm{Li}_{j}(\sqrt{-1})+\mathrm{Li}_{j}(-\sqrt{-1})\right)+(-1)^{j}\frac{(k-2)!}{(k-j)!(j-1)!}\left(\frac{\pi\sqrt{-1}}{2}\right)^{k}$ (77)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +(-1)^{k}(k-2)!\left(\mathrm{Li}_{k}(1)+\mathrm{Li}_{k}(-1)\right)-\frac{\log 2}{k-1}\left(\frac{\pi\sqrt{-1}}{2}\right)^{k-1}+\frac{1}{k-1}\left(\frac{\pi\sqrt{-1}}{2}\right)^{k}-\frac{1}{k}\left(\frac{\pi\sqrt{-1}}{2}\right)^{k}.$ (78)
@@
$\displaystyle\mathrm{I}_{g}(\omega_{2}^{k-1}\omega_{0})$ $\displaystyle=-\sum_{\begin{subarray}{c}j=2\\ j:\text{odd}\end{subarray}}^{k}(-1)^{\frac{1-j}{2}}\frac{1}{(k-j)!}\left(\frac{\pi}{2}\right)^{k-j}2^{1-j}(2^{1-j}-1)\zeta(j)$ (83)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1-(-1)^{k}}{2}(-1)^{\frac{1-k}{2}}2^{1-k}\zeta(k)+\frac{1}{(k-1)!}\log 2\left(\frac{\pi}{2}\right)^{k-1}.$ (84)
The equations 40 and 42 are shown by setting $k=2m$ and $k=2m+1$. Furthermore, by Proposition 3.3,
@@
$\displaystyle\mathrm{I}_{h}(\omega_{4}^{2m-1}\omega_{0})$ $\displaystyle=\frac{1}{3^{2m-1}}\frac{1}{(2m-1)!}\left(\frac{\pi}{\sqrt{3}}\right)^{2m-1}\frac{1}{2}\log 3$ (89)

### v3

@@ 1. Introduction
This conjecture has not yet been proved as of August 2024, but Goncharov ([7]), Terasoma ([15]), Deligne–Goncharov ([4]) showed
@@ 1.1. Main Results
Note that $X_{f}$ is defined over $\mathbb{Q}$ with good reduction outside $S_{f}$ where
$\displaystyle S_{f}=\begin{cases}\{2\},&f=g,\\ \{2,3\},&f=h.\end{cases}$ (11)
The main result of this paper is to give a motivic interpretation for these values by using motivic iterated integrals. To explain the claim, we need more notation. Fix an algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$. For a number field $F(\subseteq\overline{\mathbb{Q}})$ and the $\mathbb{Q}$-vector subspace $R\subseteq F^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ with $R=\mathcal{O}^{\times}_{F,S}\otimes_{\mathbb{Z}}\mathbb{Q}$ for some finite set $S$ of finite places of $F$, let $\mathcal{A}_{F,R}$ be the fundamental Hopf algebra of $R$ and $\mathcal{H}_{F,R}:=\mathcal{A}_{F,R}\otimes_{\mathbb{Q}}\mathbb{Q}[\tau]$. Let $\mathrm{per}:\mathcal{H}_{F,R}\rightarrow\mathbb{C}$ be the period map. Then the Galois group $\mathrm{Gal}(F/\mathbb{Q})$ acts on $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ (see Section 4 for details).
@@
where $\xi_{N}:=\exp\left(\frac{2\pi\sqrt{-1}}{N}\right)$. We define
$\displaystyle R_{f}:=\mathcal{O}_{F_{f},S^{\prime}_{f}}^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}=\begin{cases}\langle 2\rangle_{F_{g}},&f=g,\\ \langle 2,3\rangle_{F_{h}},&f=h,\end{cases}$ (12)
where
$\displaystyle S^{\prime}_{f}$ :=$\displaystyle:=\{v\text{ finite primes of }F_{f}\leavevmode\nobreak\ |\leavevmode\nobreak\ v|p\text{ for any }p\in S_{f}\}$ (13)
=$\displaystyle=\begin{cases}\{(1-\xi_{4})\},&f=g,\\ \{(2),(1-\xi_{3})\},&f=h\end{cases}$ (14)
and $\langle a_{1},\dots,a_{r}\rangle_{F}$ is the $F$-linear space spanned by $a_{1},\dots,a_{r}$. Also, we put $G_{f}:=\mathrm{Gal}(F_{f}/\mathbb{Q})$. Then, we have the following theorem.
@@
We will compute the value of $D_{f}^{(k)}$ and conjectural value of $\dim_{\mathbb{Q}}Z_{f}^{(k)}$ after Corollary 5.10 obtained by numerical experiments.
@@
We organize this paper as follows. In Section 2, we set up some notations and definitions. In Section 3, we explain some basic properties of our iterated integrals and give some explicit formulas. These formulas shows our periods can be written as linear combinations of well known periods and it leads us to seek their motivic interpretations, which are the contents in Sections 4 and 5. In Section 4, we summarize the basic properties of the motivic iterated integrals given by Deligne, Glanois, and Goncharov to give motivic interpretations of the iterated integrals on our algebraic curves. In Section 5, we give motivic interpretations and investigate their algebraic structures. This interpretation explains the appearance of $\pi$, $\log 2$, $\log 3$, $L(k,\chi_{-3})$, and Riemann zeta values in the explicit formulas in Example 3.1-(3), Propositions 3.5 and 3.7. We also notice about the direct sum decomposition of the $\mathbb{Q}$-linear space $Z_{f}^{(k)}$. In Section 5.4, we describe the relationship between our periods and special values called alternating multiple mixed values, defined by Xu–Yan–Zhao [16] and further studied by Charlton [1].
We remark that we considered in [14] only the values given by iterated integrals with three elements of $B_{g}$ and we consider a wider class in this paper.
@@ Acknowledgement
The author would like to express his deepest gratitude to his supervisor, Professor T. Yamauchi. In particular, we follow his idea in a part of the computation in Section 3 and Subsection 5.2. The author would also like to thank Professor H. Furusho of Nagoya University for inviting me to his monthly seminar and gave me various advice of my research. The author is grateful to Professor S. Charlton of the Max-Planck Institute for Mathematics for discussing alternating multiple mixed values. The author would like to thank the WISE Program for AI Electrics of Tohoku University for their financial support to carry out our study.
@@
  • For a scheme $X$ with locally of finite type over $\mathbb{C}$, let $X^{\mathrm{an}}$ be the analytification of $X$.
@@ 2. Preparations
$\displaystyle L\left(\begin{matrix}k_{1},\dots,k_{d}\\ \alpha_{1},\dots,\alpha_{d}\end{matrix}\right):=(-1)^{d}\mathrm{I}_{\mathrm{dch}}(0;\alpha_{1}^{-1},\{0\}^{k_{1}-1},\dots,\alpha_{d}^{-1},\{0\}^{k_{d}-1};1)$ (21)
@@
If $f(X,Y,Z)=X^{2}+Y^{2}-Z^{2}$, we can take a $\mathbb{Q}$-basis as $B_{f}=\left\{\frac{\leavevmode\nobreak\ dx}{x},\frac{\leavevmode\nobreak\ dx}{1-x},\frac{\leavevmode\nobreak\ dx}{y},\frac{\leavevmode\nobreak\ dx}{xy}\right\}$. Then, if we restrict the 11 11-forms to the subset $\left\{\frac{\leavevmode\nobreak\ dx}{x},\frac{\leavevmode\nobreak\ dx}{1-x},\frac{\leavevmode\nobreak\ dx}{y}\right\}\subseteq B_{f}$, the iterated integrals in the equation 9 are $F_{2}$-MZVs considered in [14].
@@
    $\displaystyle\phi_{g}:\mathbb{P}^{1}(\mathbb{C})\overset{\sim}{\longrightarrow}(X_{g})^{\mathrm{an}},\phi_{g}(\lambda):=(x,y)=\left(\frac{2\lambda}{1+\lambda^{2}},\frac{1-\lambda^{2}}{1+\lambda^{2}}\right).$ (22)
    The points on $X_{g}$ with $x=0,1,\infty$ correspond to the following five points: $(x,y)=\left(0,\pm 1\right),(1,0)$ and $[X:Y:Z]=\left[1:\pm\sqrt{-1}:0\right]$. As mentioned in [11, p.10], it holds
    $\displaystyle H^{1}_{\mathrm{dR}}(Y_{g}/\mathbb{Q})\otimes_{\mathbb{Q}}\mathbb{C}\cong H^{1}_{\mathrm{dR}}((Y_{g})^{\mathrm{an}})\cong H^{1}_{\mathrm{dR}}(\mathbb{P}^{1}(\mathbb{C})\setminus\{\text{five points}\}),$ (23)
    then the dimension of the $\mathbb{Q}$-linear space $H^{1}_{\mathrm{dR}}(Y_{g}/\mathbb{Q})$ is 44 44 and we can take $B_{g}$ as
@@
    $\displaystyle\phi_{h}:\mathbb{P}^{1}(\mathbb{C})\overset{\sim}{\longrightarrow}(X_{h})^{\mathrm{an}},\phi_{h}(\lambda):=(x,y)=\left(\frac{\lambda(\lambda+2)}{1+\lambda+\lambda^{2}},\frac{1-\lambda^{2}}{1+\lambda+\lambda^{2}}\right).$ (24)
    The points on $X_{h}$ with $x=0,1,\infty$ correspond to the following six points: $(x,y)=(0,\pm 1),(1,0),(1,-1)$ and $[X:Y:Z]=[1:\xi_{3}^{\pm 1}:0]$. Since
    $\displaystyle H^{1}_{\mathrm{dR}}(Y_{h}/\mathbb{Q})\otimes_{\mathbb{Q}}\mathbb{C}\cong H^{1}_{\mathrm{dR}}((Y_{h})^{\mathrm{an}})\cong H^{1}_{\mathrm{dR}}(\mathbb{P}^{1}(\mathbb{C})\setminus\{\text{six points}\}),$ (25)
    the dimension of the $\mathbb{Q}$-linear space $H^{1}_{\mathrm{dR}}(Y_{h}/\mathbb{Q})$ is 55 55 and we can take $B_{h}$ as
@@ 3.2. Explicit formula
where $\varphi_{j}=\sum_{n=0}^{\infty}c(n,\varphi_{j})x^{n-1}\leavevmode\nobreak\ dx$ (see [14, p.8, Proposition 3.1.3]). It implies equation 35 since $\omega_{s}=\sum_{n=0}^{\infty}c(n,\omega_{s})x^{n-1}\leavevmode\nobreak\ dx$ for each $s=0,1,\dots,6$. ∎
@@
To prove this explicit formula, we need the following lemma.
@@
If $z_{0}\in\mathbb{C}$ satisfies $|z_{0}|=1$ and $-\frac{\pi}{2}\leq\arg(z_{0})\leq\frac{\pi}{2}$, then
@@
$\displaystyle\mathrm{Li}_{k}(z_{0}):=-\mathrm{I}_{\mathrm{dch}_{0,z_{0}}}(0;1,\{0\}^{k-1};z_{0})=\sum_{n=1}^{\infty}\frac{z_{0}^{n}}{n^{k}}.$ (49)
@@
where $\gamma_{1}:=\mathrm{dch}_{0,1-\varepsilon}$, $\gamma_{2}:=C_{1}(1-\varepsilon,1+\varepsilon\sqrt{-1})$, and $\gamma_{3}:=C_{0}(1+\varepsilon\sqrt{-1},z_{2})$ for sufficiently small $\varepsilon>0$ (see FIGURE 1).
Рис. 1. integral paths
@@
By $\varepsilon\searrow 0$ and calculating the regularization of the iterated integral (Remark 2.2), then
$\displaystyle\int_{\gamma_{1}}\frac{\leavevmode\nobreak\ dz_{1}}{1-z_{1}}=0,$ (52)
@@
where $\gamma_{3}:=C_{0}(1,z_{2})$. Here,
$\displaystyle\int_{\mathrm{dch}}\frac{\leavevmode\nobreak\ dz_{1}}{1+z_{1}}=\log 2.$ (57)
@@
$\displaystyle=-\int_{-\gamma}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}$ (59)
$\displaystyle=-\int_{\gamma_{4}\gamma_{5}\gamma_{6}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z},$ (60)
where $-\gamma$ is defined by $(-\gamma)(t):=-\gamma(t)$, $\gamma_{4}:=\mathrm{dch}_{-1,-\varepsilon}$, $\gamma_{5}:=C_{0}\left(-\varepsilon,-\varepsilon\frac{z_{0}}{|z_{0}|}\right)$, and $\gamma_{6}:=\mathrm{dch}_{-\varepsilon\frac{z_{0}}{|z_{0}|},-z_{0}}$ for sufficiently small $\varepsilon>0$ (see FIGURE 2).
Рис. 2. integral paths
@@
$\displaystyle\int_{\gamma_{6}}\frac{\leavevmode\nobreak\ dz}{z}=\int_{\gamma_{5}}\frac{\leavevmode\nobreak\ dz}{1-z}=0,$ (61)
@@
$\displaystyle\int_{\gamma_{4}\gamma_{5}\gamma_{6}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}=\int_{\gamma_{4}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}+\int_{\gamma_{4}}\frac{\leavevmode\nobreak\ dz}{1-z}\int_{\gamma_{5}}\frac{\leavevmode\nobreak\ dz}{z}+\int_{\gamma_{6}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}.$ (62)
Here, each term can be calculated as
$\displaystyle\int_{\gamma_{4}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}$ $\displaystyle=\int_{\gamma_{4}^{-1}}\frac{\leavevmode\nobreak\ dz}{z}\frac{\leavevmode\nobreak\ dz}{1-z}$ (63)
@@
$\displaystyle\int_{\gamma_{4}}\frac{\leavevmode\nobreak\ dz}{1-z}\int_{\gamma_{5}}$ $\displaystyle\frac{\leavevmode\nobreak\ dz}{z}=\log 2\log z_{0},$ (66)
$\displaystyle\int_{\gamma_{6}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}$ $\displaystyle=\int_{\mathrm{dch}_{0,-z_{0}}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}=\mathrm{Li}_{2}(-z_{0}),$ (67)
@@
where $\gamma=C_{0}(1,\sqrt{-1})$. Since $0\leq\arg(z)\leq\frac{\pi}{2}$ for any $z\in\gamma([0,1])$, we have
@@
$\displaystyle=\left[\log^{k-2}z(-\mathrm{Li}_{2}(z)+\frac{\pi\sqrt{-1}}{2}\log z-\mathrm{Li}_{2}(-z)-\log 2\log z-\frac{1}{2}\log^{2}z)\right]_{1}^{\sqrt{-1}}$ (76)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -(k-2)\int_{1}^{\sqrt{-1}}\log^{k-3}z(-\mathrm{Li}_{2}(z)+\frac{\pi\sqrt{-1}}{2}\log z-\mathrm{Li}_{2}(-z)-\log 2\log z-\frac{1}{2}\log^{2}z)\frac{\leavevmode\nobreak\ dz}{z}.$ (77)
@@
$\displaystyle=\sum_{j=2}^{k}(-1)^{j-1}\frac{(k-2)!}{(k-j)!}\left\{\left(\frac{\pi\sqrt{-1}}{2}\right)^{k-j}\left(\mathrm{Li}_{j}(\sqrt{-1})+\mathrm{Li}_{j}(-\sqrt{-1})\right)\right.$ (79)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left.-\frac{j-1}{j!}\left(\frac{\pi\sqrt{-1}}{2}\right)^{k}+\frac{1}{(j-1)!}\log 2\left(\frac{\pi\sqrt{-1}}{2}\right)^{k-1}\right\}$ (80)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +(-1)^{k}(k-2)!\left(\mathrm{Li}_{k}(1)+\mathrm{Li}_{k}(-1)\right).$ (81)
@@
$\displaystyle\mathrm{I}_{g}(\omega_{2}^{k-1}\omega_{0})$ $\displaystyle=-\frac{1}{(k-2)!}\left\{\sum_{\begin{subarray}{c}j=2\\ j:\text{odd}\end{subarray}}^{k}(-1)^{\frac{1-j}{2}}\frac{(k-2)!}{(k-j)!}\left(\frac{\pi}{2}\right)^{k-j}2^{1-j}(2^{1-j}-1)\zeta(j)\right.$ (84)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left.+\sum_{j=2}^{k}\frac{(k-2)!}{(k-j)!(j-1)!}(-1)^{j-1}\log 2\left(\frac{\pi}{2}\right)^{k-1}\right.$ (85)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left.+\frac{1-(-1)^{k}}{2}(-1)^{\frac{1-k}{2}}(k-2)!2^{1-k}\zeta(k)\right\}$ (86)
$\displaystyle=-\sum_{\begin{subarray}{c}j=2\\ j:\text{odd}\end{subarray}}^{k}(-1)^{\frac{1-j}{2}}\frac{1}{(k-j)!}\left(\frac{\pi}{2}\right)^{k-j}2^{1-j}(2^{1-j}-1)\zeta(j)$ (87)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{(k-1)!}\log 2\left(\frac{\pi}{2}\right)^{k-1}+\frac{1-(-1)^{k}}{2}(-1)^{\frac{1-k}{2}}2^{1-k}\zeta(k).$ (88)
The equations 42 and 44 are shown by setting $k=2m$ and $k=2m+1$. Furthermore, by Proposition 3.3,
@@
$\displaystyle\mathrm{I}_{h}(\omega_{4}^{2m-1}\omega_{0})$ $\displaystyle=\frac{1}{\sqrt{3}^{2m-1}}\frac{1}{(2m-1)!}\left(\frac{\pi}{\sqrt{3}}\right)^{2m-1}\frac{1}{2}\log 3$ (93)